A stuntman is designing a movie scene. The actor will “fall” from a radio tower. He needs to know how fast the actor will be moving after falling from the 30m tower to provide the proper airbag for safety. How fast will he be falling?


1
Expert's answer
2021-04-23T07:27:43-0400

"\begin{cases}
v=gt,\\
s=\frac{gt^2}{2};
\end{cases}" "\implies" "\begin{cases}
t=\sqrt{\frac{2s}{g}},\\
v=\sqrt{2gs};
\end{cases}"

"t=2.5~s,"

"v=24~\frac ms."